OpenStudy (anonymous):

the square root of "x+7" +5=x please solve step by step. Thanks

OpenStudy (anonymous):

\[\sqrt{x+7}+5=x\]

OpenStudy (campbell_st):

\[\sqrt{x+7} + 5 = x\] square every term \[x + 7 + 25 = x^2\] rewriting \[x^2 - x -32 = 0\]

OpenStudy (campbell_st):

solve the quadratic by GQF

OpenStudy (anonymous):

dont you have to isolate the radical first??

OpenStudy (campbell_st):

you are solving for x... lots of ways of doing it

OpenStudy (anonymous):

i thought i subtract 5 from both sides then square each side.

OpenStudy (anonymous):

\[\begin{align} \sqrt{x+7}+5&=x\\ \sqrt{x+7}&=x-5\\ x+7&=(x-5)^2\\ x+7&=x^2-10x+25\\ 0&=x^2-11x+18\\ 0&=(x-9)(x-2)\\ x&=9,2 \end{align}\]

OpenStudy (anonymous):

Campbell, you squared incorrectly. You can't just square the 5, if you leave the 5 on the other side you have to square it as a binomial.

OpenStudy (anonymous):

Thanks!!

OpenStudy (campbell_st):

your suggestion \[\sqrt{x+7}=x-5\] \[x + 7 = x^2 - 10x + 25\]

OpenStudy (lgbasallote):

\((\sqrt{x-7} + 5)^2\) is not squaring each term @campbell_st :D

OpenStudy (anonymous):

nbouscal can u help me with 1 more?

OpenStudy (anonymous):

Close this question and start a new one if you have another question :)

OpenStudy (anonymous):

okay dont wanna lose you lol.

OpenStudy (anonymous):

I'm not goin' anywhere :P